What is the voltmeter reading across the R2 resistor in a series circuit containing a 12 V potential and three resistors R1= 3 ohms, R2=2 ohms, and R3= 1 ohm?


1
Expert's answer
2021-05-09T15:16:56-0400

We have given the Circuit Potential "V = 12V"


Three resistance are connected in series,


then equivalent resistance = "2+3+1 = 6\Omega"


Current in the circuit, "I = \dfrac{12}{6} = 2A"


Hence, Voltmeter reading across resistance "R_2 = IR_2 = 2 \times 2 = 4V"